# Neurosymbolic Grounding for Compositional World Models

### v1

@@ 3.1 Slot-based Autoencoder
An autoencoder transforms an image into a hidden representation and then reconstructs the image from this representation. We use a slot-based autoencoder to learn a factorized hidden representation, where each factor represents a single entity. Such autoencoder make two assumptions: (1) each factor captures a specific property of the image, and (2) collectively, all factors describe the entire input image. This set-structured representation enables unsupervised disentanglement of objects into individual entities. Our slot based autoencoder consists of two components:
$\textsc{EntityExtractor}:I\rightarrow\{S_{1},\dots,S_{k}\}$: The entity extractor takes an image and produces a set-structured hidden representation describing each object. [23] introduce an iterative-refining based attention algorithm to learn such an encoding (slot attention). In our domains, slot attention and derivate works [5, 18] struggle to disentangle images into separate slots. To avoid the perception model becoming a bottleneck for studying dynamics learning, we propose a new entity extractor that uses Segment Anything [22] with pretrained weights to produce set-structured segmentation masks to decompose the image into objects and a Resnet-18 image encoder [15] to produce a set-structured hidden representation for each object. This model allows us to perfectly match ground truth segmentations in our domains while preserving the assumptions of set structured hidden representations.
$\textsc{SpatialDecoder}:\{S_{1},\dots,S_{k}\}\rightarrow\{I_{1}^{\prime},I_{2}^{\prime},I_{k}^{\prime}\},\{M_{1},M_{2},M_{k}\}$: The slot decoder is a permutation equivariant network that decodes the set of slots back to the input image. We follow previous works [23, 44, 9] in using a spatial decoder [41] that decodes a given set of vectors $\{S_{1},\dots,S_{k}\}$ individually into a set of reconstructions $\{I_{1}^{\prime},I_{2}^{\prime},I_{k}^{\prime}\}$ and a set of masks $\{M_{1},\dots,M_{k}\}$. The final image $I$ is produced by taking the Hadamard product of each reconstruction and its corresponding mask and adding all the resulting images together. That is, $I=\sum_{i=1}^{k}M_{i}\cdot I_{i}^{\prime}$.
@@ 3.2 Neurosymbolic encoding
To achieve robust CG in our setting, our representation must be resilient to both object and attribute compositions. The $k$ slot-based encoding, by construction, generalizes to object replacement. For attribute compositions, however, it is essential to know the exact attributes to be targeted for CG. Given these attributes, we propose describing each entity with a composition of symbol vectors. Each symbol vector is associated with a single entity and attribute, allowing it to trivially generalize to different attributes compositions. Moreover, we can ensure a canonical ordering of the symbols, making downstream attention-based computations invariant to permutations of attributes. We’ll next detail our method for generating these symbol vectors.
@@
$\textsc{Slm}:I_{i}\times\{C_{1},C_{2},C_{p}\}\rightarrow\Lambda_{i}:=(\Lambda_{i}^{C_{1}};\dots;\Lambda_{i}^{C_{p}})$: The symbolic labelling module (Slm) processes an image and a predefined list of attributes. Assuming this list is comprehensive (though not exhaustive), the module employs a pretrained Clip model to compute attention scores between the image features and each entity encoding. The resulting logits indicate the alignment between the image and each attribute. The attribute most aligned with the image is then identified using a straight-through Gumbel softmax [17]. The gumbel-softmax yields the index of the most likely value for each attribute as one-hot vectors, which are concatenated together to form a bit-vector. However, the discrete representation of such bit-vectors do not align well with downstream attention based modules so, instead of directly using one-hot vectors, the gumbel-softmax selects a value-specific encoding for each attribute. Thus, the resultant symbol vector is a composition of learnable latent encodings distinct to each attribute value. In implementation, Slm utilizes another zero-shot symbolic module (a spatial-softmax operation [41]) to extract positional attributes, such as the $x$ and $y$ position of the object, from the disentangled input vector.
3.2.1 Action Conditioning
@@
$\textsc{ActionAttn}:\{\Lambda_{1},\Lambda_{2},\Lambda_{k}\}\times\{A_{1},A_{2},A_{k}\}\rightarrow\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\}$: To predict the next state accurately, world models need to condition the state on the intended action. This conditioning is achieved by concatenating the action to the encoding. Ideally, if each entity occupies one slot and the slots have a canonical ordering, this conditioning can be achieved by simply concatenating the corresponding action to the corresponding encoding. However, the entities in a set structured representation do not have a fixed order. To find such a canonical ordering, we follow [44] in learning a permutation matrix between the actions and the slots and using the permutation matrix to reorder the slots and concatenate them with their respective actions. Attention, by construction, is equivariant with respect to slot order, which avoids the need to enforce a canonical ordering.
@@ 3.3 Transition Model
Monolithic transition models like GNNs and MLPs model every pairwise object interaction, which, in domains with sparse object interactions, leads to accumulating spurious correlations. Our modular transition model addresses this problem by selecting a relevant pairwise interaction and updating the encodings for those objects locally (following [9]). We model this selection process by computing key-query neurosymbolic attention between ordered symbolic and neural rule encodings to determine the most applicable rule-slot pair. As each entity is composed of shared symbol vectors, the dot-product activations between the symbolic encodings and the rule encodings remain consistent for objects with identical attributes. Next, we will discuss the mechanisms of module selection and transitions.
$\textsc{ModuleSelector}:\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\}\times\{\vec{R}_{1},\vec{R}_{2},\vec{R}_{l}\}\rightarrow(p,c,r)$: The goal of the selection process is to select the primary slot, the contextual slot (which the primary slot will be conditioned on), and the update function to model the interaction. Query-key attention [3] serves as the base mechanism we will be using to perform selections. We compute query-key attention between the rule encoding and the action-conditioned symbolic encoding. The naive algorithm to compute this selection will take $O(k^{2}l)$ time, where $k$ is the number of slots and $l$ is the number of rules. In implementation, the selection of $(r,p,c)$ can be reduced to a runtime of $O(kl+k)$ by partial application of the query-key attention. This algorithm is presented in appendix section 2.
@@ 4 Experiments
Figure 3: A single update step of Cosmos. The image $I$ is fed through a slot-based autoencoder and a CLIP model to generate the slot encodings $\{S_{1},S_{2},S_{k}\}$ and the symbol vectors $\{\Lambda_{1},\Lambda_{2},\Lambda_{k}\}$. The actions and the symbolic encoding are aligned and concatenated using a permutation equivariant action attention module, which are used to select the update rule to be applied to the slots. This figure depicts a single update step; in implementation, the update-select-transform step is repeated multiple times to model multi-object interactions.
@@
We demonstrate the effectiveness of Cosmos on the 2D Block pushing domain [21, 44, 9, 19] with scene composition, and two instances of object composition. We selected the 2D block pushing domain as it is a widely-adopted and well-studied benchmark in the community. Notably, even in this synthetic domain, our instances of compositional generalization proved challenging to surpass for baseline models, underscoring the difficulty of the problem.
@@
2D Block Pushing Domain: The 2D block pushing domain [21, 19] is a two-dimensional environment that necessitates dynamic and perceptual reasoning. Figure 5 presents an overview of this domain. All objects have four attributes: color ($\Lambda^{C_{\texttt{color}}}$), shape ($\Lambda^{C_{\texttt{shape}}}$), $x$ position ($\Lambda^{C_{\texttt{x-pos}}}$), and $y$ position ($\Lambda^{C_{\texttt{y-pos}}}$). Objects can be pushed in one of the four cardinal directions (North-East-South-West). Heavier objects can push lighter objects, but not the other way around. The weight of the object depends on the shape attribute. At each step, the agent observes an image of size $3\times 224\times 224$ with $k$ objects and an action pushing one of the objects. This action is chosen from a uniform random distribution. The goal is for the agent to capture the dynamics of object movement. Furthermore, while there are $n=|A_{\texttt{color}}|\times|A_{\texttt{shape}}|$ unique objects in total, only $k<n$ objects are allowed to co-occur in a realized scene.
@@
Dataset Setup: We adapt the methodology from [21, 44] with minor changes. For scene compositions (SC), we construct training and testing to have unique object combinations. In object composition (OC), objects with matching attributes exhibit identical dynamics. Two specific cases are explored: Team Composition (OC-Team) where dynamics are shared based on color, and Sticky Composition (OC-Sticky) where dynamics are shared based on color and adjacency. Further details are in the appendix (Section A.1). Our data generation methodology ensures that the compound distribution is disjoint, while the atom distribution remains consistent across datasets, i.e. $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{train})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})=\emptyset$ and $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{train})=\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{eval})$. The difficulty of the domain can be raised by increasing the number of objects. We sample datasets for 3 and 5 objects.
@@
Evaluation: We compare against past works in compositional world modeling with publically available codebases at the time of writing. For the block pushing domain, we compare against Homomorphic world models (Gnn) [44] and NPS [9] with modifications to ensure that the actions are aligned with the slots (AlignedNps). The latter is equivalent to an ablation of Cosmos without the symbolic labelling module. For compositional world modeling, we analyzed next-state predictions using mean squared error (MSE), current-state reconstructions using auto-encoder mean squared error (AE-MSE), and latent state separation using (MRR) following previous work [21, 44, 9]. Recognizing limitations in existing MRR calculation, we introduce the Equivariant MRR (Eq.MRR) metric, which accounts for different slot orders when calculating the MRR score. More details can be found in the appendix Section A.3.
@@
3 objects 5 objects Dataset Model MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ OC Cosmos 4.23E-03 4.90E-04 1.20E-01 4.15E-03 1.68E-03 3.67E-01 (Sticky) AlignedNps 1.14E-02 7.72E-03 8.01E-02 6.07E-03 2.47E-03 3.62E-01 Gnn 7.94E-03 5.11E-03 6.03E-04 6.21E-03 2.73E-03 5.30E-04 OC Cosmos 4.60E-03 4.33E-04 1.04E-01 5.53E-03 1.86E-03 2.86E-01 (Team) AlignedNps 1.24E-02 8.36E-03 1.75E-01 9.64E-03 3.12E-03 2.93E-01 Gnn 8.92E-03 3.82E-03 7.16E-04 7.01E-03 1.62E-03 5.46E-04 SC Cosmos 7.66E-04 6.34E-05 2.99E-01 4.08E-04 2.92E-06 3.03E-01 AlignedNps 3.51E-03 2.69E-03 2.97E-01 2.45E-03 1.22E-03 3.19E-01 Gnn 9.89E-03 1.03E-02 5.50E-01 1.20E-02 1.28E-02 5.25E-01
@@
Table 1: Evaluation results on the 2D block pushing domain for scene composition (SC) and object composition (OC) averaged across three seeds. We report next-state reconstruction error (MSE), autoencoder reconstruction error (AE-MSE), and the equivariant mean reciprocal rank (Eq.MRR) for three transition models: our model (Cosmos), an improved version of [9] (AlignedNps), and a reimplementation of [44] (Gnn). Our model (Cosmos) achieves best next-state reconstructions for all datasets.
@@
Results: Results are presented in Table 2. First, we find that Cosmos achieves the best next state prediction performance (MSE) on all benchmarks. Moving from scene composition to object composition datasets shows a drop in performance, underscoring the complexity of the task. Surprisingly, there is less than expected performance degradation moving from the three object to five object domain. We attribute this to the nature of the block pushing domain and the choice of loss function. MSE loss measures the pixel error between the predicted and next image. As the density of objects in the image increases, the reconstruction target becomes more informative, encouraging better self-correction, and hence more efficient training.
@@
Finally, we do not notice a consistent pattern in the Equivariant MRR scores between models. First, all models tend to exhibit a higher Eq.MRR score in the five object environments. However, in many cases, models with high Eq.MRR score also have underperforming autoencoders. For instance, in the five object scene composition case, the GNN exhibits a high Eq.MRR score yet simultaneously has the worst autoencoder reconstruction error. We notice this happens when the model suffers a partial representation collapse (overfitting to certain color, shapes combination seen during training). This maps many slot encodings to the same neighborhood in latent space; making it easier to retrieve similar encodings, boosting the MRR score. Given these observations, we conclude that MRR might not be an optimal indicator of a model’s downstream utility. For a comprehensive assessment of downstream utility, we turn to the methodology outlined by [36], applying the models to a downstream planning task.
Downstream Utility: We evaluate the downstream utility of all world models using a simple planning task in 2D shapeworld environments. Our method, inspired by [36], involves using a greedy planner that acts based on the Hungarian distance between the predicted and goal states. Due to the compounding nature of actions over a trajectory, there is an observed divergence from the ground truth the deeper we get into the trajectory. We run these experiments on our test dataset and average the scores at each trajectory depth. We showcase results in Figure 4. Our model (Cosmos) shows the most consistency and least deviation from the goal state in all datasets, which suggests that neurosymbolic grounding helps improve the downstream efficacy of world models. More details can be found in appendix section A.4.

### v2

@@ 3.1 Slot-based Autoencoder
A slot-based autoencoder transforms an image into a factorized hidden representation, with each factor representing a single entity, and then reconstructs the image from this representation. Such autoencoder make two assumptions: (1) each factor captures a specific property of the image (2) collectively, all factors describe the entire input image. This set-structured representation enables unsupervised disentanglement of objects into individual entities. Our slot based autoencoder has two components:
$\textsc{EntityExtractor}:I\rightarrow\{S_{1},\dots,S_{k}\}$: The entity extractor takes an image and produces a set-structured hidden representation describing each object. In our domains, slot attention and derivate works (Chang et al., 2022; Jia et al., 2022) struggle to disentangle images into separate slots. To avoid the perception model becoming a bottleneck for studying dynamics learning, we propose a new entity extractor that uses Segment Anything Kirillov et al. (2023) with pretrained weights to produce set-structured segmentation masks to decompose the image into objects and a Resnet-18 image encoder He et al. (2016) to produce a set-structured hidden representation for each object. This model allows us to perfectly match ground truth segmentations in our domains while preserving the assumptions of set structured hidden representations. More details are presented in Section A.1.
$\textsc{SpatialDecoder}:\{S_{1},\dots,S_{k}\}\rightarrow\{I_{1}^{\prime},\dots I_{k}^{\prime}\},\{M_{1}^{\prime},\dots M_{k}^{\prime}\}$: The slot decoder is a permutation equivariant network that decodes the set of slots back to the input image. We follow previous works (Locatello et al., 2020; Zhao et al., 2022; Goyal et al., 2021) in using a spatial decoder (Watters et al., 2019b) that decodes a given set of vectors $\{S_{1},\dots,S_{k}\}$ individually into a set of image reconstructions $\{I_{1}^{\prime},\dots I_{k}^{\prime}\}$ and a set of mask reconstructions $\{M_{1}^{\prime},\dots,M_{k}^{\prime}\}$. The final image $I$ is produced by taking the Hadamard product of each reconstruction and its corresponding mask and adding all the resulting images together. That is, $I=\sum_{i=1}^{k}M_{i}^{\prime}\cdot I_{i}^{\prime}$.
@@ 3.2 Neurosymbolic encoding
To achieve robust CompGen in our setting, our representation must be resilient to both object and attribute compositions. The $k$ slot-based encoding, by construction, generalizes to object replacement. For attribute compositions, however, it is essential to know the exact attributes to be targeted for CompGen. Given these attributes, we propose describing each entity with a composition of symbol vectors. Each symbol vector is associated with a single entity and attribute, allowing it to trivially generalize to different attributes compositions. Moreover, we can ensure a canonical ordering of the symbols, making downstream attention-based computations invariant to permutations of attributes. We’ll next detail our method for generating these symbol vectors.
@@
$\textsc{Slm}:I_{i}\times\{C_{1},\dots C_{p}\}\rightarrow\Lambda_{i}:=(\Lambda_{i}^{C_{1}};\dots;\Lambda_{i}^{C_{p}})$: The symbolic labelling module (Slm) processes an image and a predefined list of attributes. Assuming this list is comprehensive (though not exhaustive), the module employs a pretrained Clip model to compute attention scores between the image features and each entity encoding. The resulting logits indicate the alignment between the image and each attribute. The attribute most aligned with the image is then identified using a straight-through Gumbel softmax (Jang et al., 2016). The gumbel-softmax yields the index of the most likely value for each attribute as one-hot vectors, which are concatenated together to form a bit-vector. However, the discrete representation of such bit-vectors do not align well with downstream attention based modules so, instead of directly using one-hot vectors, the gumbel-softmax selects a value-specific encoding for each attribute. Thus, the resultant symbol vector is a composition of learnable latent encodings distinct to each attribute value. In implementation, Slm utilizes another zero-shot symbolic module (a spatial-softmax operation (Watters et al., 2019b)) to extract positional attributes, such as the $x$ and $y$ position of the object, from the disentangled input vector.
For instance, in Figure 3, Slm takes in the slot corresponding to the circle and a list of attributes (shape, color, etc.) and selects the most relevant attribute value (‘circle’, ‘red’, etc.). We subsequently select trainable embeddings corresponding to each attribute value and concatenate them to construct the symbolic embedding.
@@
$\textsc{ActionAttn}:\{\Lambda_{1},\dots\Lambda_{k}\}\times\{A_{1},\dots A_{k}\}\rightarrow\{\overline{\Lambda}_{1},\dots,\overline{\Lambda}_{k}\}$: For accurate next state prediction, world models must condition the state on the corresponding action. Typically, this is done by concatenating each action to its associated encoding if the slots have a canonical ordering. However, the entities in a set structured representation do not have a fixed order. To find such a canonical ordering, we follow Zhao et al. (2022) in learning a permutation matrix between the actions and the slots and using the permutation matrix to reorder the slots and concatenate them with their respective actions. Attention, by construction, is equivariant with respect to slot order, which avoids the need to enforce a canonical ordering. For instance, in Figure 3, if $\Lambda_{1}$ corresponds to the circle, a trained ActionAttn module reorders the actions so that $A_{1}$ corresponds to going downwards.
@@ 3.3 Transition Model
Monolithic transition models like GNNs and MLPs model every pairwise object interaction, leading to spurious correlations in domains with sparse interactions. Our modular transition model addresses this problem by selecting a relevant pairwise interaction and updating the encodings for those objects locally (following Goyal et al. (2021)). We model this selection process by computing key-query neurosymbolic attention between ordered symbolic and neural rule encodings to determine the most applicable rule-slot pair. As each entity is composed of shared symbol vectors, the dot-product activations between the symbolic encodings and the rule encodings remain consistent for objects with identical attributes. Next, we will discuss the mechanisms of module selection and transitions.
$\textsc{ModuleSelector}:\{\overline{\Lambda}_{1},\dots\overline{\Lambda}_{k}\}\times\{\vec{R}_{1},\dots\vec{R}_{l}\}\rightarrow(p,c,r)$: The goal of the selection process is to select the primary slot, the contextual slot (which the primary slot will be conditioned on), and the update function to model the interaction. Query-key attention (Bahdanau et al., 2015) serves as the base mechanism we will be using to perform selections. We compute query-key attention between the rule encoding and the action-conditioned symbolic encoding. The naive algorithm to compute this selection will take $O(k^{2}l)$ time, where $k$ is the number of slots and $l$ is the number of rules. In implementation, the selection of $(p,c,r)$ can be reduced to a runtime of $O(kl+k)$ by partial application of the query-key attention. This algorithm is presented in Section 2 of the appendix.
@@ 4 Experiments
Figure 3: A single update step of Cosmos. The image $I$ is fed through a slot-based autoencoder and a CLIP model to generate the slot encodings $\{S_{1},\dots S_{k}\}$ and the symbol vectors $\{\Lambda_{1},\dots\Lambda_{k}\}$. The actions and the symbolic encoding are aligned and concatenated using a permutation equivariant action attention module, which are used to select the update rule to be applied to the slots. This figure depicts a single update step; in implementation, the update-select-transform step is repeated multiple times to model multi-object interactions.
@@
We demonstrate the effectiveness of Cosmos on the 2D Block pushing domain (Kipf et al., 2019; Zhao et al., 2022; Goyal et al., 2021; Ke et al., 2021) with entity composition, and two instances of relational composition. We selected the 2D block pushing domain as it is a widely-adopted and well-studied benchmark in the community. Notably, even in this synthetic domain, our instances of compositional generalization proved challenging to surpass for baseline models, underscoring the difficulty of the problem.
@@
2D Block Pushing Domain: The 2D block pushing domain (Kipf et al., 2019; Ke et al., 2021) is a two-dimensional environment that necessitates dynamic and perceptual reasoning. Figure 5 presents an overview of this domain. All objects have four attributes: color ($\Lambda^{C_{\texttt{color}}}$), shape ($\Lambda^{C_{\texttt{shape}}}$), $x$ position ($\Lambda^{C_{\texttt{x-pos}}}$), and $y$ position ($\Lambda^{C_{\texttt{y-pos}}}$). Objects can be pushed in one of the four cardinal directions (North-East-South-West). Heavier objects can push lighter objects, but not the other way around. The weight of the object depends on the shape attribute. At each step, the agent observes an image of size $3\times 224\times 224$ with $k$ objects and an action pushing one of the objects. This action is chosen from a uniform random distribution. The goal is for the agent to capture the dynamics of object movement. Furthermore, while there are $n=|A_{\texttt{color}}|\times|A_{\texttt{shape}}|$ unique objects in total, only $k<n$ objects are allowed to co-occur in a realized scene.
@@
Dataset Setup: We adapt the methodology from (Kipf et al., 2019; Zhao et al., 2022) with minor changes. For entity compositions (EC), we construct training and testing datasets to have unique object combinations between them. In relational composition (RC), objects with matching attributes exhibit identical dynamics. Two specific cases are explored: Team Composition (RC-Team) where dynamics are shared based on color, and Sticky Composition (RC-Sticky) where dynamics are shared based on color and adjacency. Further details are in the appendix (Section A.2). Our data generation methodology ensures that the compound distribution is disjoint, while the atom distribution remains consistent across datasets, i.e. $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{train})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})=\emptyset$ and $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{train})=\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{eval})$. The difficulty of the domain can be raised by increasing the number of objects. We sample datasets for 3 and 5 objects.
@@
Evaluation: We compare against past works in compositional world modeling with publically available codebases at the time of writing. For the block pushing domain, we compare with homomorphic world models (Gnn) (Zhao et al., 2022) and NPS (Goyal et al., 2021) with modifications to ensure that the actions are aligned with the slots (AlignedNps). The latter is equivalent to an ablation of Cosmos without the symbolic labelling module. We analyzed next-state predictions using mean squared error (MSE), current-state reconstructions using auto-encoder mean squared error (AE-MSE), and latent state separation using (MRR) following previous work (Kipf et al., 2019; Zhao et al., 2022; Goyal et al., 2021). Recognizing limitations in existing MRR calculation, we introduce the Equivariant MRR (Eq.MRR) metric, which accounts for different slot orders when calculating the MRR score. More details can be found in the appendix Section A.4.
@@
3 objects 5 objects Dataset Model MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ RC Cosmos 4.23E-03 4.90E-04 1.20E-01 4.15E-03 1.68E-03 3.67E-01 (Sticky) AlignedNps 1.14E-02 7.72E-03 8.01E-02 6.07E-03 2.47E-03 3.62E-01 Gnn 7.94E-03 5.11E-03 6.03E-04 6.21E-03 2.73E-03 5.30E-04 RC Cosmos 4.60E-03 4.33E-04 1.04E-01 5.53E-03 1.86E-03 2.86E-01 (Team) AlignedNps 1.24E-02 8.36E-03 1.75E-01 9.64E-03 3.12E-03 2.93E-01 Gnn 8.92E-03 3.82E-03 7.16E-04 7.01E-03 1.62E-03 5.46E-04 EC Cosmos 7.66E-04 6.34E-05 2.99E-01 4.08E-04 2.92E-06 3.03E-01 AlignedNps 3.51E-03 2.69E-03 2.97E-01 2.45E-03 1.22E-03 3.19E-01 Gnn 9.89E-03 1.03E-02 5.50E-01 1.20E-02 1.28E-02 5.25E-01
@@
Table 1: Evaluation results on the 2D block pushing domain for entity composition (EC) and relational composition (RC) averaged across three seeds. We report next-state reconstruction error (MSE), autoencoder reconstruction error (AE-MSE), and the equivariant mean reciprocal rank (Eq.MRR) for three transition models: our model (Cosmos), an improved version of Goyal et al. (2021) (AlignedNps), and a reimplementation of Zhao et al. (2022) (Gnn). Our model (Cosmos) achieves best next-state reconstructions for all datasets.
@@
Results: Results are presented in Table 2. First, we find that Cosmos achieves the best next state prediction performance (MSE) on all benchmarks. Moving from entity composition to relational composition datasets shows a drop in performance, underscoring the complexity of the task. Surprisingly, there is less than expected performance degradation moving from the three object to five object domain. We attribute this to the nature of the block pushing domain and the choice of loss function. MSE loss measures the pixel error between the predicted and next image. As the density of objects in the image increases, the reconstruction target becomes more informative, encouraging better self-correction, and hence more efficient training.
@@
Finally, we do not notice a consistent pattern in the Equivariant MRR scores between models. First, all models tend to exhibit a higher Eq.MRR score in the five object environments. However, in many cases, models with high Eq.MRR score also have underperforming autoencoders. For instance, in the five object entity composition case, the GNN exhibits a high Eq.MRR score yet simultaneously has the worst autoencoder reconstruction error. We notice this happens when the model suffers a partial representation collapse (overfitting to certain color, shapes combination seen during training). This maps many slot encodings to the same neighborhood in latent space; making it easier to retrieve similar encodings, boosting the MRR score. Given these observations, we conclude that MRR might not be an optimal indicator of a model’s downstream utility. For a comprehensive assessment of downstream utility, we turn to the methodology outlined by Veerapaneni et al. (2020), applying the models to a downstream planning task.
@@
Downstream Utility: We evaluate the downstream utility of all world models using a simple planning task in 2D shapeworld environments. Our method, inspired by Veerapaneni et al. (2020), involves using a greedy planner that acts based on the Hungarian distance between the predicted and goal states. Due to the compounding nature of actions over a trajectory, there is an observed divergence from the ground truth the deeper we get into the trajectory. We run these experiments on our test dataset and average the scores at each trajectory depth. We showcase results in Figure 4. Our model (Cosmos) shows the most consistency and least deviation from the goal state in all datasets, which suggests that neurosymbolic grounding helps improve the downstream efficacy of world models. More details can be found in appendix section A.5.